# t-DGR: A Trajectory-Based Deep Generative Replay Method for Continual Learning in Decision Making

### v1

@@ 5.2 Metrics
We evaluate our models using three metrics proposed by the Continual World benchmark [8], with the average success rate being the primary metric. Although the forward transfer and forgetting metrics are not well-defined in a “General Continual Learning" setting, they are informative within the context of Continual World benchmarks. As a reminder from Section 3.2, let $N$ denote the number of tasks, and $s_{i}$ represent the success rate of the learner on task $\mathcal{T}_{i}$. Additionally, let $s_{i}(t)$ denote the success rate of the learner on task $\mathcal{T}_{i}$ after training on tasks $\mathcal{T}_{1}$ to $\mathcal{T}_{t}$.
@@ 5.3 Baselines
  • oEWC [18]: A variation of EWC known as online Elastic Weight Consolidation (oEWC) bounds the memory of EWC by employing a single penalty term for the previous model instead of individual penalty terms for each task. This baseline is the representative regularization-based method.
  • PackNet [19]: This baseline is the representative parameter isolation method. Packnet safeguards previous task information in a neural network by iteratively pruning, freezing, and retraining parts of the network.
  • DGR [7]: This baseline is a deep generative replay method that only generates individual state observations i.i.d. and not entire trajectories.
  • CRIL [29]: This baseline is a deep generative replay method that trains a policy along with a start state generator and a dynamics model that predicts the next state given the current state and action. Trajectories are generated by using the dynamics model and policy to autoregressively generate next states from a start state.
@@
Due to the inability of oEWC and PackNet to handle blurry task boundaries, we made several adjustments for CW20 and GCL10. Since PackNet cannot continue training parameters for a task once they have been fixed, we treated the second repetition of tasks in CW20 as distinct from the first iteration, resulting in PackNet being evaluated with $N=20$, while the other methods were evaluated with $N=10$. As for GCL10 and its blurry task boundaries, the best approach we could adopt with oEWC and PackNet was to apply their regularization techniques at regular training intervals rather than strictly at task boundaries. During evaluation, all tasks were assessed using the last fixed set of parameters in the case of PackNet.
@@
DGR 75.0 $\pm$5.8 -4.3 $\pm$5.1 17.8 $\pm$4.1
CRIL 28.4 $\pm$10.6 -1.1 $\pm$2.8 68.6 $\pm$10.4
t-DGR 81.9 $\pm$3.3 -0.3 $\pm$4.9 14.4 $\pm$2.5
(b) GCL10
@@
DGR 75.3 $\pm$4.4
CRIL 53.5 $\pm$5.5
t-DGR 81.7 $\pm$4.0
@@
DGR 74.1 $\pm$4.1 18.9 $\pm$2.9 23.3 $\pm$3.3
CRIL 50.8 $\pm$4.4 4.4 $\pm$4.9 46.1 $\pm$5.4
t-DGR 83.9 $\pm$3.0 30.6 $\pm$4.5 14.6 $\pm$2.9
@@
Table 1: Tables (a), (b), and (c) present the results for Continual World 10, General Continual Learning 10, and Continual World 20, respectively. The tables display the average success rate, forward transfer, and forgetting (if applicable) with 90% confidence intervals using 5 random seeds. An up arrow indicates that higher values are better and a down arrow indicates that smaller values are better. Table (d) compares the impact of replay amount on the average success rate of t-DGR and DGR on CW10 with 90% confidence intervals obtained using 5 random seeds. The best results are highlighted in bold.
@@ 5.4 Discussion
t-DGR emerges as the leading method, demonstrating the highest success rate on CW10 (Table 0(a)), CW20 (Table 0(c)), and GCL10 (Table 0(b)). Notably, PackNet’s performance on the second iteration of tasks in CW20 diminishes, highlighting its limited capacity for continually accommodating new tasks. This limitation underscores the fact that PackNet falls short of being a true lifelong learner, as it necessitates prior knowledge of the task count for appropriate parameter capacity allocation. On the contrary, pseudo-rehearsal methods, such as t-DGR, exhibit improved performance with the second iteration of tasks in CW20 due to an increased replay time. These findings emphasize the ability of DGR methods to effectively leverage past knowledge, as evidenced by their superior forward transfer in both CW10 and CW20.
@@
GCL10 (Table 0(b)) demonstrates that pseudo-rehearsal methods are mostly unaffected by blurry task boundaries, whereas PackNet’s success rate experiences a significant drop-off. This discrepancy arises from the fact that PackNet’s regularization technique does not work effectively with less clearly defined task boundaries.
@@
Additionally, the diminishing performance gap between DGR and t-DGR as the replay ratio increases in Table 0(d) indicates that a higher replay ratio reduces the likelihood of any portion of the trajectory being insufficiently covered when sampling individual state observations i.i.d., thereby contributing to improved performance. This trend supports the theoretical sample complexity of DGR derived in Section 2.2, as $\Theta(n\log n+mn\log\log n)$ closely approximates the sample complexity of t-DGR, $\Theta(mn)$, when the replay amount $m\to\infty$. However, while DGR can achieve comparable performance to t-DGR with a high replay ratio, the availability of extensive replay time is often limited in many real-world applications.
@@ 6 Conclusion
Importantly, t-DGR aligns with the concept of “General Continual Learning" by taking into account essential properties of the real world, including bounded memory and blurry task boundaries. These considerations ensure that our method remains applicable and effective in real-world scenarios, enabling its potential integration into practical applications.
@@ Appendix B Model Architecture
Table 2: Multi-layer perceptron architecture of the learner shared by all methods

### v2

@@ 5.2 Metrics
We evaluate our models using three metrics proposed by the Continual World benchmark (Wolczyk et al., 2021), with the average success rate being the primary metric. Although the forward transfer and forgetting metrics are not well-defined in a setting with blurry task boundaries, they are informative within the context of Continual World benchmarks. As a reminder from Section 3.2, let $N$ denote the number of tasks, and $s_{i}$ represent the success rate of the learner on task $\mathcal{T}_{i}$. Additionally, let $s_{i}(t)$ denote the success rate of the learner on task $\mathcal{T}_{i}$ after training on tasks $\mathcal{T}_{1}$ to $\mathcal{T}_{t}$.
@@ 5.3 Baselines
  • oEWC (Schwarz et al., 2018): A variation of EWC known as online Elastic Weight Consolidation (oEWC) bounds the memory of EWC by employing a single penalty term for the previous model instead of individual penalty terms for each task. This baseline is the representative regularization-based method.
  • PackNet (Mallya & Lazebnik, 2018): This baseline is the representative parameter isolation method. Packnet safeguards previous task information in a neural network by iteratively pruning, freezing, and retraining parts of the network.
  • DGR (Shin et al., 2017): This baseline is a deep generative replay method that only generates individual state observations i.i.d. and not entire trajectories.
  • CRIL (Gao et al., 2021): This baseline is a deep generative replay method that trains a policy along with a start state generator and a dynamics model that predicts the next state given the current state and action. Trajectories are generated by using the dynamics model and policy to autoregressively generate next states from a start state.
@@
Due to the inability of oEWC and PackNet to handle blurry task boundaries, we made several adjustments for CW20 and BB10. Since PackNet cannot continue training parameters for a task once they have been fixed, we treated the second repetition of tasks in CW20 as distinct from the first iteration, resulting in PackNet being evaluated with $N=20$, while the other methods were evaluated with $N=10$. As for BB10 and its blurry task boundaries, the best approach we could adopt with oEWC and PackNet was to apply their regularization techniques at regular training intervals rather than strictly at task boundaries. During evaluation, all tasks were assessed using the last fixed set of parameters in the case of PackNet.
5.4 Ablations
To evaluate the effect of generator architecture on pseudo-rehearsal methods, we included an ablation where the diffusion models in pseudo-rehearsal methods are replaced with Wasserstein generative adversarial networks (GAN) with gradient penalty (Gulrajani et al., 2017). To evaluate the effectiveness of diffusion models at generative replay, we evaluated the quality of generated samples for past tasks at all stages of continual learning. Since we utilize proprioceptive observations rather than RGB camera observations, there is no practical way to qualitatively evaluate the generated samples. The proprioceptive observations are 39-dimensional, comprising 3D positions and quaternions of objects, and the degree to which the robot gripper is open (Yu et al., 2020). These raw numbers cannot be qualitatively evaluated like RGB images. Instead, we use the average diffusion loss as a quantitative proxy metric for generation quality. When the generator model is learning task $i$, we compute the average diffusion loss for data samples from tasks 1 to $i-1$. We then compare this average diffusion loss to the average diffusion loss when the generator model first learned task $i$. The difference between the two provides an estimate of how much generative replay has degraded the generation quality of previous tasks.
@@ 5.4 Ablations
DGR (gan) 21.2 $\pm$3.7 -2.6 $\pm$3.3 74.4 $\pm$4.7
CRIL (gan) 24.0 $\pm$4.4 0.3 $\pm$4.5 72.0 $\pm$5.4
t-DGR (gan) 17.4 $\pm$4.4 0.1 $\pm$3.1 79.8 $\pm$4.2
DGR (diffusion) 75.0 $\pm$5.8 -4.3 $\pm$5.1 17.8 $\pm$4.1
CRIL (diffusion) 28.4 $\pm$10.6 -1.1 $\pm$2.8 68.6 $\pm$10.4
t-DGR (diffusion) 81.9 $\pm$3.3 -0.3 $\pm$4.9 14.4 $\pm$2.5
(b) BB10
@@
DGR (gan) 35.6 $\pm$4.3
CRIL (gan) 37.5 $\pm$4.7
t-DGR (gan) 28.3 $\pm$3.9
DGR (diffusion) 75.3 $\pm$4.4
CRIL (diffusion) 53.5 $\pm$5.5
t-DGR (diffusion) 81.7 $\pm$4.0
@@
DGR (gan) 19.0 $\pm$2.1 0.8 $\pm$2.3 78.6 $\pm$2.3
CRIL (gan) 26.7 $\pm$6.5 -0.6 $\pm$0.7 72.1 $\pm$6.2
t-DGR (gan) 20.6 $\pm$6.1 0.7 $\pm$3.4 76.3 $\pm$5.3
DGR (diffusion) 74.1 $\pm$4.1 18.9 $\pm$2.9 23.3 $\pm$3.3
CRIL (diffusion) 50.8 $\pm$4.4 4.4 $\pm$4.9 46.1 $\pm$5.4
t-DGR (diffusion) 83.9 $\pm$3.0 30.6 $\pm$4.5 14.6 $\pm$2.9
@@
Table 1: Tables (a), (b), and (c) present the results for Continual World 10, Blurry Boundaries 10, and Continual World 20, respectively. The tables display the average success rate, forward transfer, and forgetting (if applicable) with 90% confidence intervals using 5 random seeds. An up arrow indicates that higher values are better and a down arrow indicates that smaller values are better. Table (d) compares the impact of replay amount on the average success rate of t-DGR and DGR on CW10 with 90% confidence intervals obtained using 5 random seeds. The best results are highlighted in bold.
@@ 5.5 Discussion
t-DGR emerges as the leading method, demonstrating the highest success rate on CW10 (Table 1(a)), CW20 (Table 1(c)), and BB10 (Table 1(b)). Notably, PackNet’s performance on the second iteration of tasks in CW20 diminishes, highlighting its limited capacity for continually accommodating new tasks. This limitation underscores the fact that PackNet falls short of being a true lifelong learner, as it necessitates prior knowledge of the task count for appropriate parameter capacity allocation. On the contrary, pseudo-rehearsal methods, such as t-DGR, exhibit improved performance with the second iteration of tasks in CW20 due to an increased replay time. These findings emphasize the ability of DGR methods to effectively leverage past knowledge, as evidenced by their superior forward transfer in both CW10 and CW20.
@@
BB10 (Table 1(b)) demonstrates that pseudo-rehearsal methods are mostly unaffected by blurry task boundaries, whereas PackNet’s success rate experiences a significant drop-off. This discrepancy arises from the fact that PackNet’s regularization technique does not work effectively with less clearly defined task boundaries.
In our experiments across CW10, CW20, and BB10, we observed that diffusion models outperform GANs as the generator for pseudo-rehearsal methods. We hypothesize that the distributional shifts in continual learning exacerbate instability issues with GAN training (Salimans et al., 2016; Brock et al., 2018; Miyato et al., 2018). The motivation behind CRIL as stated in the paper (Gao et al., 2021) is to alleviate the burden of trajectory generation from the generator by transferring some of the generation complexity to a dynamics model. Our findings support this motivation when working with less capable generators (Dhariwal & Nichol, 2021), as reducing the generator’s burden consistently enhances performance of GAN-based pseudo-rehearsal methods across all benchmarks. Among pseudo-rehearsal techniques, CRIL places the least demand on its generator, which only needs to produce the initial frame of a trajectory. In contrast, DGR requires generation of every frame, and t-DGR, the most demanding, requires generation of every frame along with the handling of trajectory timestep conditioning. Results from our GAN-based pseudo-rehearsal experiments on CW10, CW20, and BB10 indicate that CRIL outperforms DGR, which in turn outperforms t-DGR. However, experiments using a more capable diffusion generator (Dhariwal & Nichol, 2021) reveal a reversal in performance ranking among these methods, suggesting that more capable generators diminish the need to offload generation complexity from the main generator.
@@
The diminishing performance gap between DGR and t-DGR as the replay ratio increases in Table 1(d) indicates that a higher replay ratio reduces the likelihood of any portion of the trajectory being insufficiently covered when sampling individual state observations i.i.d., thereby contributing to improved performance. This trend supports the theoretical sample complexity of DGR derived in Section 2.2, as $\Theta(n\log n+mn\log\log n)$ closely approximates the sample complexity of t-DGR, $\Theta(mn)$, when the replay amount $m\to\infty$. However, while DGR can achieve comparable performance to t-DGR with a high replay ratio, the availability of extensive replay time is often limited in many real-world applications.
Recent studies have indicated that applying generative replay to diffusion models for image data leads to a collapse of generation quality due to compounding errors in the generated synthetic data (Zajac et al., 2023; Masip et al., 2023; Smith et al., 2023). However, our experiments reveal that when generating lower dimensional proprioceptive data, the diffusion model is able to maintain generation quality. Although Figure 3 shows that generative replay degrades generation quality of past tasks as the diffusion model learns new tasks, the success rates in Table 1 suggest that this degradation in generation quality is not severe enough to impact the learner’s performance. Notably, generation quality for past tasks appears to deteriorate in tasks 5 and 7 but improves in the subsequent task, suggesting that the error compounding in diffusion models during generative replay might not be as severe as previously assumed, particularly for lower-dimensional data.
@@
Figure 3: This table illustrates the ability of the diffusion model in t-DGR to generate past data as it continues to learn additional tasks in CW10 through generative replay. The line plot for task $i$ plots the average diffusion loss of the diffusion model in future tasks on task $i$ data. The loss is an L1 version of the diffusion training loss in Equation 6.
@@ 6 Conclusion
Importantly, t-DGR takes into account essential properties of the real world, including bounded memory and blurry task boundaries. These considerations ensure that our method remains applicable and effective in real-world scenarios, enabling its potential integration into practical applications.
@@ B.1 Multi-layer Perceptron
Table 2: Multi-layer perceptron architecture of the learner shared by all methods, featuring five linear layers with ReLU activations applied after each layer except the final one.
To condition the MLP learner on the task, the input vector is concatenated with a one-hot vector representing the task before being passed into the MLP.
B.2 Diffusion U-net
The U-net consists of 4 downsampling and 4 upsampling layers. To condition the generator U-net on the trajectory timestep, we use a sinusoidal positional encoder from Transformers to create an sinusoidal positional embedding that gets passed through 3 linear layers and then added to the convolutional layer output at each level of the U-net.
B.3 Generative Adversarial Network
The generator is implemented with an MLP consisting of 7 linear layers with batch normalization and Leaky ReLU activations applied after each layer except the final one.
The discriminator is implemented with an MLP consisting of 3 linear layers with Leaky ReLU activations applied after each layer except the final one.
Sinusoidoal positional encodings are fed as input to a 2-layer MLP to condition both the generator and discriminator on the trajectory timestep for t-DGR.